What Use the intercept form to find the equation of the line with the given intercepts. X-intercept (-1/6,0) y-intercept (0,-2/3)
Answer by Alan3354(69443)

Use the intercept form to find the equation of the line with the given intercepts. X-intercept (-1/6,0) y-intercept (0,-2/3)
--------
I left out the "What"
-----
If the intercepts are (A,0) andn (0,B), the equation is
Bx + Ay = A*B
-2x/3 - y/6 = 1/9
12x + 3y = -2